OpenStudy (anonymous):

be sure to simplify by collectiong like terms before you solve 8x-5x=15

OpenStudy (anonymous):

x=

OpenStudy (anonymous):

5

OpenStudy (anonymous):

\(8x-5x=15 \implies 3x=15 \implies x=5\).

OpenStudy (anonymous):

x = 5. 8x - 5x = 3x. 15 divided by 3 = 5. x = 5

OpenStudy (anonymous):

8x -5x = x(8-5) = 3x 3x = 15 which means x = 5

OpenStudy (anonymous):

you should know the distributive, associative and one other property by now. I forget the last one :)

OpenStudy (anonymous):

Zero.

OpenStudy (anonymous):

It's called the Zero Propert,

OpenStudy (anonymous):

identity you mean.

OpenStudy (anonymous):

Whatever.

OpenStudy (anonymous):

oh you are so cool!